# TheAgentCompany: Benchmarking LLM Agents on Consequential Real World Tasks

### v1

@@ Abstract
We interact with computers on an everyday basis, be it in everyday life or work, and many aspects of work can be done entirely with access to a computer and the Internet. At the same time, thanks to improvements in large language models (LLMs), there has also been a rapid development in AI agents that interact with and affect change in their surrounding environments. But how performant are AI agents at helping to accelerate or even autonomously perform work-related tasks? The answer to this question has important implications for both industry looking to adopt AI into their workflows, and for economic policy to understand the effects that adoption of AI may have on the labor market. To measure the progress of these LLM agents’ performance on performing real-world professional tasks, in this paper, we introduce TheAgentCompany, an extensible benchmark for evaluating AI agents that interact with the world in similar ways to those of a digital worker: by browsing the Web, writing code, running programs, and communicating with other coworkers. We build a self-contained environment with internal web sites and data that mimics a small software company environment, and create a variety of tasks that may be performed by workers in such a company. We test baseline agents powered by both closed API-based and open-weights language models (LMs), and find that with the most competitive agent, 24% of the tasks can be completed autonomously. This paints a nuanced picture on task automation with LM agents – in a setting simulating a real workplace, a good portion of simpler tasks could be solved autonomously, but more difficult long-horizon tasks are still beyond the reach of current systems.
@@ 1 Introduction
We are in the midst of a technological transformation. With the rapid year-by-year and month-by-month progress brought about by large language models (LLMs), we are seeing AI-based assistance or automation become commonplace in tasks that were unthinkable only a few years ago. In fact, the pace of progress is so fast that some have gone so far as to claim that the majority of human labor may be automatable within the next couple of years (Eloundou et al., 2023; Amodei & Fridman, 2024). On the other hand, others are skeptical, claiming that language models cannot truly reason (Kambhampati et al., 2024), do not generalize well to novel tasks (Chollet et al., 2024), and may only have an impact on a small minority of the labor market (Wittenstein, 2024).
@@
We perform experiments using seven large language model backbones, including API-based models such as Anthropic Claude (Anthropic, 2023), OpenAI GPT-4o (OpenAI, 2024), Google Gemini (Team et al., 2023), Amazon Nova (Intelligence, 2024), as well as open models including Meta Llama (Dubey et al., 2024) and Alibaba Qwen (Yang et al., 2024). All models are run using the OpenHands agent framework (Wang et al., 2024b), [^1] https://github.com/All-Hands-AI/OpenHands which provides a stable and strong agent harness for both web browsing and coding. As a result of experiments, we find that the best performing model, Claude 3.5 Sonnet was able to autonomously perform 24.0% of the provided tests to completion, and achieve a score of 34.4% on our metric that provides extra credit for partially completed tasks.
@@
In the rest of this paper, we explain detail comparisons to other existing benchmarks (§ 2), how we set up realistic and reproducible environments (§ 3), how we define tasks (§ 4) and how we create them (§ 5), our baseline agent (§ 6), experimental results (§ 7), and finally implications and future directions (§ 8).
Table 1: Comparison of different AI agent benchmarks. Interface: the interface agent has access to; [Uncaptioned image] is web browser, [Uncaptioned image] is desktop, [Uncaptioned image] is API usage, [Uncaptioned image] is Python script, [Uncaptioned image] is chat platform, [Uncaptioned image] is bash terminal. Supported Tasks: tasks in the benchmark, * indicate tasks with no association with real-world occupations; SE refers to software engineering, HR is human resources, PM is project manager. Checkpoint-based evaluation: if tasks are evaluated at intermediate checkpoints and assigned partial scores. Interact with NPC Agents: If the agent can interact with other NPC agents during task-solving.
Framework Diverse Real-world Work Task Categories Requires Interaction Long-Horizon w/ Checkpoints Interface Self-Hosted Environment MiniWob++ (Liu et al., 2018) Browsing [Uncaptioned image] Mind2Web (Deng et al., 2023) Browsing [Uncaptioned image] WebLINX (Lù et al., 2024) Browsing [Uncaptioned image] AssistantBench (Yoran et al., 2024) Browsing [Uncaptioned image] WebArena (Zhou et al., 2023) Browsing [Uncaptioned image] VisualWebArena (Koh et al., 2024) Browsing [Uncaptioned image] VideoWebArena (Jang et al., 2024) Browsing [Uncaptioned image] WorkArena (Drouin et al., 2024) Enterprise Software [Uncaptioned image] OSWorld (Xie et al., 2024) Office, Coding [Uncaptioned image] [Uncaptioned image] Windows Agent Arena (Bonatti et al., 2024) Browsing, Office, Coding [Uncaptioned image] [Uncaptioned image] AppWorld (Trivedi et al., 2024) Daily [Uncaptioned image] Gorilla APIBench (Patil et al., 2023) Coding [Uncaptioned image] $\tau$-bench (Yao et al., 2024) Retail, Airline [Uncaptioned image] SWE-bench (Jimenez et al., 2024) SWE [Uncaptioned image] DevBench (Li et al., 2024) SWE [Uncaptioned image] Smallville (Park et al., 2023) Social [Uncaptioned image] Sotopia (Zhou et al., 2024) Social [Uncaptioned image] TheAgentCompany SWE, HR, Admin, PM, Research, Finance [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 2 Benchmark Desiderata and Comparison to Other Benchmarks
In order to evaluate the ability of agents to perform tasks in complex real-world settings, we built TheAgentCompany with a number of desiderata in mind. The comparison with several existing prominent agent benchmarks with respect to these desiderata is in Table 1.
Coverage of Multiple Work-related Tasks:
In order to make any valid statements about the potential of AI to accelerate or automate various types of real-world work, we should have tasks that are motivated by real-world work across multiple job categories. Many benchmarks are not relevant to real-world work (e.g. MiniWob++ (Liu et al., 2018)) or very relevant to real-world work, but only over a limited scope of tasks (e.g. SWE-Bench (Jimenez et al., 2024)). In contrast, TheAgentCompany contains a set of more diverse, realistic, and professional tasks that would typically be completed by multiple job roles in a software engineering company.
Requirement for Interaction:
If agents are to integrate into real-world workplaces, they will need to communicate with the other human members of the workspace. Most other benchmarks do not measure communication or interactivity, with the exception of $\tau$-bench (Yao et al., 2024), which only measures interaction in customer service scenarios. TheAgentCompany provides a better testbed for communication as many tasks involve asking and providing information to colleagues as part of a more complex task.
Long-horizon Tasks with Checkpoints:
In real-world settings, many tasks require taking many different steps to achieve a higher-level goal. One major novel contribution of TheAgentCompany is that we both (1) contain tasks that require an agent to perform significantly more consecutive work (i.e. involving more steps and realistically taking human professionals longer to accomplish) than previous benchmarks, and (2) provide granular evaluators that measure the ability of models to perform subtasks of these larger tasks.
Versatile Environment Interface:
In order to handle a diversity of tasks in real-world settings, we minimally should be able to interact with the tools that real-world workers use – including web interfaces, programs, command-line terminals, and communication tools. TheAgentCompany covers all of these interfaces, while most previous benchmarks focus only on one or two.
Self-hosted and Reproducible:
In order to allow for careful comparisons between different methods that remain constant over time, the benchmark should be fully self-hosted and reproducible. This contrasts with existing benchmarks that do not have execution environments (e.g. Mind2Web (Deng et al., 2023)) or require the usage of third-party software (e.g. WorkArena (Drouin et al., 2024)).
@@ 3 TheAgentCompany Environment Setup
Our benchmark is set in an imaginary software engineering startup called TheAgentCompany, hence the benchmark’s name. Within TheAgentCompany, we create tasks inspired by tasks handled by workers inside such companies. More details about the company’s imaginary background, overview and employees can be found in Appendix A. The benchmark environment contains multiple components.
Local Workspace
The local workspace runs locally on the agent’s host, which is analogous to a human professional’s local workspace, e.g. their work laptop computer. This environment is created as a sandboxed Docker environment to provide a safe execution environment that will not affect other parts of the evaluation machine (Wang et al., 2024b). [^2] https://docs.all-hands.dev/modules/usage/how-to/custom-sandbox-guide This environment is where agents work on the task, and within this environment the TheAgentCompany baseline agent (§ 6) uses a browser, code editor and a Linux terminal with typical software preinstalled.333 Other options would include using something like a GUI-based desktop environment with office software (Xie et al., 2024), but we opt to build a baseline solution that is entirely web-based, reflecting the recent trend of more enterprise software moving to the cloud.
Intranet
This part of the environment mimics the company’s internal websites that host code, documents, project management software, and communications software. To achieve our goal of a reproducible, self-contained environment, we follow WebArena (Zhou et al., 2023), in using open-source, self-hostable software to host our environment. The environment mainly contains the following websites:
@@
Simulated Colleague Communication
One major aspect of working in a company is communicating with other company members, and in TheAgentCompany we also test the ability of models to perform this type of communication. Specifically, we allow agents to use RocketChat to message other company members and obtain information that may not be available in the original task description. To create these simulated colleagues, we rely on the Sotopia platform (Zhou et al., 2024), which supports the creation of simulated human characters with LLMs. Each simulated colleague is equipped with a detailed profile that includes their name, role, responsibilities, and project affiliations. (e.g., Sarah Johnson, who serves as the CTO, oversees technical strategy planning and R&D team leadership, with access to all technical channels). Agents can interact with these simulated colleagues through direct messages or in specific channels, as is standard in RocketChat and other platforms. By default, all simulated human characters are backed by the Claude-3-5-Sonnet-20241022 LLM across experiments, as we found that it provided the best results during preliminary experiments. For example conversations between the agent and the simulated colleagues drawn from empirical experiments, please refer to Appendix B.
@@ 4 Task Structure
The tasks in TheAgentCompany include a task intent, a list of checkpoints that the agent must achieve, a programmatic evaluator to check success on these checkpoints, and code to initialize and finalize the environment. We show some examples in Table 2, and describe each of aspect in detail below.
Table 2: Example task intents and checkpoints for three domains.
Task Intent
Each task begins with an English description, simulating how a user would instruct an LLM-based agent to perform a real-world task. In general, we aim for these tasks to be clear enough so that a human worker would be able to complete the task without asking for further instructions directly from the user (although they may need to ask questions of their other co-workers).
Checkpoints
Tasks are divided into checkpoints representing intermediate milestones, each assigned a point value to measure progress. Each checkpoint is awarded a certain number of points based on its significance to the overall completion of the task. Checkpoints are written in English, and typically specify one or more of the following:
@@
Evaluators
Checkpoints are created in the task design phase, but for actual evaluation, each of the checkpoints must be concretely implemented through an evaluator – a program that checks the completion of the checkpoint. These evaluators are implemented by examining environment states, such as the local workspace, intranet status, simulated colleague interactions, or by analyzing agent trajectories, like verifying browsing history or action sequences.
In most cases, these evaluators are deterministic and written as simple Python functions. For instance, in the SWE task in Table 2, the checkpoints are deterministic: verifying if the JanusGraph repository is cloned, the binary file is built, and the server is launched with an HTTP endpoint. However, for tasks with more complex and unstructured deliverables, such as in Table 2, the last checkpoint in the Finance task requires contacting the correct finance director (David Wong) to resolve ambiguous questions, which involves a judgment from a (simulated) human colleague, deterministic evaluation can be challenging due to subjectivity and variability. In such cases, we employ LLM-based evaluation. This involves prompting LLMs with predefined rubrics or reference outputs to assess the agent’s deliverables, enabling a more nuanced and flexible evaluation of these tasks. Same as the NPC backbone, all LLM-based evaluators are backed by the Claude-3-5-Sonnet-20241022.
@@ 4.1 Evaluation Metrics
Full completion score
We define the full completion score $S_{\text{full}}$ as:
@@
This binary metric evaluates whether the agent successfully completed the task by passing all checkpoints.
Partial completion score
To provide a more nuanced measure that rewards partial task completion while strongly incentivizing full task completion, we define partial completion score $S_{\text{partial}}$ as:
$S_{\text{partial}}=0.5\cdot\frac{\text{Result}}{\text{Total}}+0.5\cdot S_{\text{full}},$
where:
  • Result: Sum of awarded points across all checkpoints (including partial credit),
  • Total: Sum of the total points for all checkpoints,
  • $\frac{\text{Result}}{\text{Total}}$: Fractional progress toward full completion,
  • $S_{\text{full}}$: Binary indicator equal to $1$ when the task is fully completed.
@@
Number of steps
The number of steps is defined as the total number of LLM calls made during the task execution. This metric quantifies the operational effort required to perform the task.
Cost per instance
The cost per instance measures the monetary cost of querying the underlying LLM through its API to complete a task. Assuming no prompt caching, the cost is calculated as:
@@
This efficiency metric reflects the computational expense of task completion based on token usage.
@@ 4.2 Workflow
Each task typically follows a workflow involving the following stages:
  1. 1.
    Initialization: The agent sets up its workspace and prepares to execute the task.
  2. 2.

### v2

@@ Abstract
We interact with computers on an everyday basis, be it in everyday life or work, and many aspects of work can be done entirely with access to a computer and the Internet. At the same time, thanks to improvements in large language models (LLMs), there has also been a rapid development in AI agents that interact with and affect change in their surrounding environments. But how performant are AI agents at accelerating or even autonomously performing work-related tasks? The answer to this question has important implications both for industry looking to adopt AI into their workflows and for economic policy to understand the effects that adoption of AI may have on the labor market. To measure the progress of these LLM agents’ performance on performing real-world professional tasks, in this paper we introduce TheAgentCompany, an extensible benchmark for evaluating AI agents that interact with the world in similar ways to those of a digital worker: by browsing the Web, writing code, running programs, and communicating with other coworkers. We build a self-contained environment with internal web sites and data that mimics a small software company environment, and create a variety of tasks that may be performed by workers in such a company. We test baseline agents powered by both closed API-based and open-weights language models (LMs), and find that the most competitive agent can complete 30% of tasks autonomously. This paints a nuanced picture on task automation with LM agents–in a setting simulating a real workplace, a good portion of simpler tasks could be solved autonomously, but more difficult long-horizon tasks are still beyond the reach of current systems. We release code, data, environment, and experiments on https://the-agent-company.com.
@@ 1 Introduction
We are in the midst of a technological transformation. With the rapid month-by-month progress brought about by large language models (LLMs), we are seeing AI-based assistance or automation become commonplace in tasks that were unthinkable only years ago. In fact, the pace of progress is so fast that some have gone so far as to claim that the majority of human labor may be automatable within the next couple of years (Eloundou et al., 2023; Amodei & Fridman, 2024). On the other hand, others are skeptical, claiming that language models cannot truly reason (Kambhampati et al., 2024), do not generalize well to novel tasks (Chollet et al., 2024), and may only have an impact on a small minority of the labor market (Wittenstein, 2024).
@@
We perform experiments using twelve large language model backbones, including closed models such as Anthropic Claude (Anthropic, 2023), OpenAI GPT-4o (OpenAI, 2024), Google Gemini (Team et al., 2023), Amazon Nova (Intelligence, 2024), plus open models like Meta Llama (Dubey et al., 2024) and Alibaba Qwen (Yang et al., 2024). All models are run with OpenHands agent framework (Wang et al., 2024b), [^1] https://github.com/All-Hands-AI/OpenHands which provides a stable and strong agent harness for both web browsing and coding. We find in experiments that the best performing model, Gemini 2.5 Pro was able to autonomously perform 30.3% of the provided tests to completion, and achieve a score of 39.3% on our metric that provides extra credit for partially completed tasks.
@@ 2 Benchmark Desiderata and Comparison to Other Benchmarks
In order to evaluate the ability of agents to perform tasks in complex real-world settings, we built TheAgentCompany with a number of desiderata in mind. The comparison with several existing prominent agent benchmarks with respect to these desiderata is in Table 2. More details of the benchmark construction can be found in Appendix B.
Coverage of Multiple Work-related Tasks:  In order to make any valid statements about the potential of AI to accelerate or automate various types of real-world work, we should have tasks that are motivated by real-world work across multiple job categories. Many benchmarks are not relevant to real-world work (e.g. MiniWob++ (Liu et al., 2018)) or very relevant to real-world work, but only over a limited scope of tasks (e.g. SWE-Bench (Jimenez et al., 2024)). In contrast, TheAgentCompany contains a set of more diverse, realistic, and professional tasks that would typically be completed by multiple job roles in a software engineering company.
Requirement for Interaction  If agents are integrated into real-world workplaces, they need to communicate with the other human members of the workspace. Most other benchmarks do not measure communication or interactivity, except for $\tau$-bench (Yao et al., 2024) that only measures interaction in customer service scenarios. TheAgentCompany is a better testbed for communication, as asking and providing information to colleagues as part of many more complex tasks.
Long-horizon Tasks with Checkpoints  In real-world settings, many tasks require many steps to achieve a higher-level goal. One novel contribution of TheAgentCompany is that we both (1) contain tasks that require an agent to perform significantly more consecutive work (i.e. involving more steps and realistically taking human professionals longer to accomplish) than previous benchmarks, and (2) provide granular evaluators that measure the ability of models to perform subtasks of larger tasks.
Versatile Environment Interface:  In order to handle a diversity of tasks in real-world settings, we minimally should be able to interact with the tools that real-world workers use – including web interfaces, programs, command-line terminals, and communication tools. TheAgentCompany covers all of these interfaces, while most previous benchmarks focus only on one or two.
Self-hosted and Reproducible:  In order to allow for careful comparisons between different methods that remain constant over time, the benchmark should be fully self-hosted and reproducible. This contrasts with existing benchmarks that do not have execution environments (e.g. Mind2Web (Deng et al., 2023)) or require the usage of third-party hosted platform (e.g. WorkArena (Drouin et al., 2024), CRMArena (Huang et al., 2024)).
@@ 3 TheAgentCompany Environment Setup
Our benchmark is set in an imaginary software engineering startup called TheAgentCompany, hence the benchmark’s name. We create tasks inspired by tasks handled by workers inside such companies. More details about the company’s imaginary background, overview and employees can be found in Appendix G. The benchmark environment contains multiple components.
Local Workspace  The local workspace runs locally on the agent’s host, which is analogous to a human professional’s local workspace, e.g. their work laptop computer. This environment is created as a sandboxed Docker environment to provide a safe execution environment that will not affect other parts of the evaluation machine. This environment is where agents work on the task, and within this environment the TheAgentCompany baseline agent (§ 6) uses a browser, code editor and a Linux terminal with typical software preinstalled.222 Other options would include using a GUI-based desktop environment with office software (Xie et al., 2024), but we opt to build a baseline solution that is entirely web-based, reflecting the recent trend of more enterprise software moving to the cloud. Despite this, we also provide a virtual machine OS image with the entire environment pre-packaged.
Intranet  This part of the environment mimics the company’s internal websites that host code, documents, project management software, and communications software. To achieve a reproducible, self-contained environment, we follow WebArena (Zhou et al., 2023), in using open-source, self-hostable software to host our environment. The environment mainly contains the following websites:
@@
Simulated Colleague Communication  One major aspect of working in a company is communicating with other company members, and in TheAgentCompany we also test the ability of models to perform this type of communication. Specifically, we allow agents to use RocketChat to message other company members and obtain information that may not be available in the original task description. To create these simulated colleagues, we rely on the Sotopia platform (Zhou et al., 2024), which supports the creation of simulated human characters with LLMs. Each simulated colleague is equipped with a detailed profile that includes their name, role, responsibilities, and project affiliations (e.g., Sarah Johnson, who serves as the CTO, oversees technical strategy planning and R&D team leadership, with access to all technical channels). Agents can interact with these simulated colleagues through direct messages or in specific channels, as is standard in RocketChat and other platforms. By default, all simulated human characters are backed by the Claude-3-5-Sonnet-20241022 LLM across experiments, as we found that it provided the best results during preliminary experiments. The detailed error analysis with respect to the introduction of LLM as Colleageus is given in Appendix F. For example conversations between the agent and the simulated colleagues drawn from empirical experiments, please refer to Appendix H.
@@ 4 Task Structure
The tasks in TheAgentCompany include a task intent, a list of checkpoints the agent must achieve, a programmatic evaluator to check success on these checkpoints, and code to initialize and finalize the environment. We show some examples in Appendix C (Table 3), and detail each aspect below.
Task Intent  Each task begins with an English description, simulating how a user would instruct an LLM-based agent to perform a real-world task. In general, we aim for these tasks to be clear enough so that a human worker would be able to complete the task without asking for further instructions directly from the user (although they may need to ask questions of their other co-workers).
Checkpoints  Tasks are divided into checkpoints representing intermediate milestones, each assigned a point value to measure progress. Each checkpoint is awarded a certain number of points based on its significance to the overall completion of the task. Checkpoints are written in English, and typically specify one or more of the following:
@@
Evaluators  Checkpoints are created in the task design phase, but for actual evaluation, each of the checkpoints must be concretely implemented through an evaluator – a program that checks the completion of the checkpoint. These evaluators are implemented by examining environment states, such as the local workspace, intranet status, simulated colleague interactions, or by analyzing agent trajectories, like verifying browsing history or action sequences.
In most cases, these evaluators are deterministic and written as simple Python functions. For instance, in the SWE task in Table 3, the checkpoints are deterministic: verifying if the JanusGraph repository is cloned, the binary file is built, and the server is launched with an HTTP endpoint. However, for tasks with more complex and unstructured deliverables, such as in Table 3, the last checkpoint in the Finance task requires contacting the correct finance director (David Wong) to resolve ambiguous questions, which involves a judgment from a (simulated) human colleague, deterministic evaluation can be challenging due to subjectivity and variability. In such cases, we employ LLM-based evaluation. This involves prompting LLMs with predefined rubrics or reference outputs to assess the agent’s deliverables, enabling a more nuanced and flexible evaluation of these tasks. Same as the NPC backbone, all LLM-based evaluators are backed by the Claude-3-5-Sonnet-20241022. For an error analysis with respect to the LLM evaluator, refer to Appendix E.
@@ 4.1 Evaluation Metrics
Full completion score  We define the full completion score $S_{\text{full}}$ as:
@@
This binary metric evaluates if the agent successfully completed the task by passing all checkpoints.
Partial completion score  To provide a more nuanced measure that rewards partial task completion while strongly incentivizing full task completion, we define partial completion score as: $S_{\text{partial}}=0.5\cdot\frac{\text{Result}}{\text{Total}}+0.5\cdot S_{\text{full}}$, where: “Result” is sum of awarded points across all checkpoints (including partial credit), “Total” is sum of the total points for all checkpoints, $\frac{\text{Result}}{\text{Total}}$ is fractional progress toward full completion, and $S_{\text{full}}$ is binary indicator equal to $1$ when the task is fully completed.
@@
Number of steps  The number of steps is defined as the total number of LLM calls made during the task execution. This metric quantifies the operational effort required to perform the task.